# Generative Modeling with Bayesian Sample Inference

### v2

@@ 6.1 ImageNet32
For this evaluation, we train a DiT-L-2 (Peebles & Xie, 2023) in the BFN, VDM and BSI model, respectively, on the official 32×32 version of ImageNet (Chrabaszcz et al., 2017). We train each model from three seeds for $2\text{\,}\mathrm{M}$ steps and evaluate the log-likelihood of the test set in bits per dimension (BPD) and the sample quality in terms of Fréchet inception distance (FID) against the test set. For the log-likelihood, we evaluate each model’s ELBO with $5$ samples from the respective equivalent of $\mathcal{L}_{\mathrm{M}}^{\infty}$ and $2$ samples from the respective equivalent of $\mathcal{L}_{\mathrm{R}}$. For the sample quality, we draw 50 00050\,000 unconditional samples from each model with $1024$ steps and then compute the FID between the generated samples and the test set.
Table˜2 shows that BSI achieves equivalent likelihoods to VDM and BFN while generating higher-quality samples in terms of FID. This aligns with the result for consistency models by Song & Dhariwal (2024) that a larger variance of the initial state – initial belief ${\bm{\mu}}_{0}$ for BSI – improves the sample diversity. Ordering the models by improving FID, we have BFN first with an initial variance of 0 (${\bm{\mu}}_{0}={\bm{0}}$), then VDM with an initial variance of $1$ and finally BSI with an initial variance of $\lambda_{0}^{\shortminus{}\hskip-0.61455pt1}=100$. Furthermore, Fig.˜3 shows that BSI generates significantly better samples than the closely related BFN with fewer steps.
@@ 6.2 CIFAR10
Model Training Steps BPD
VDM $10\text{\,}\mathrm{M}$ 2.65
BSI 2.64
BFN $5\text{\,}\mathrm{M}$ 2.66
BSI 2.65
We train the same U-Net architecture as VDM (Kingma et al., 2023) and BFN (Graves et al., 2023) on CIFAR10. Table˜3 shows that BSI achieves equivalent log-likelihoods in terms of BPD. Due to the significant number of training steps ($10\text{\,}\mathrm{M}$), we followed (Kingma et al., 2023; Graves et al., 2023) and trained only a single model on this dataset.
@@ Acknowledgments
This research was funded by the Bavarian State Ministry for Science and the Arts within the framework of the Geothermal Alliance Bavaria project. We are thankful to David Lüdke and Jonas Dornbusch for their valuable feedback.
Software
For our results, we rely on excellent software packages, notably numpy (Harris et al., 2020), pytorch (Paszke et al., 2019), einops (Rogozhnikov, 2022), matplotlib (Hunter, 2007), h5py (Collette, 2013), hydra (Yadan, 2019) and jupyter (Granger & Pérez, 2021).
Licenses
The CIFAR10 dataset111www.cs.toronto.edu/ kriz/cifar.html has been released without a license. Access to the ImageNet dataset222image-net.org is granted under a custom license that allows non-commercial research such as this work.
@@ A.2 Diffusion Models
While DMs initially achieved prominence in image generation (Dhariwal & Nichol, 2021), they have since been applied successfully across a variety of domains, such as text-to-image mapping (Saharia et al., 2022), fluid simulations (Lienen et al., 2024; Saydemir et al., 2024), adversarial attacks (Kollovieh et al., 2024a), temporal (Lüdke et al., 2023) and general point processes (Lüdke et al., 2024), molecular dynamics (Lewis et al., 2024), molecular structure generation (Ayadi et al., 2024), and time series forecasting (Kollovieh et al., 2023, 2024b).
@@
As we have shown in Section˜A.1, BFN are a special case of our generative framework in Section˜3 if we choose $\gamma_{0}=\infty$. Curiously, Eq.˜24 shows that this is the only prior on ${\bm{\mu}}_{0}$ for which the associated forward process is Markov as the coefficient of $\bm{x}$} becomes 0. This agrees with Xue et al. (2024), who have shown that BFN admit a formulation based on stochastic differential equations, like score-based DMs.
@@ Appendix C Preconditioning Derivation
up to constant factors only depending on $\lambda$. If we plug in Eq.˜36 and isolate $f_{{\bm{\theta}}}^{\prime}$, this distance becomes
@@
Finally, we can plug $c_{\mathrm{skip}}$ into Eq.˜44 to get
@@ Appendix D Proofs
we can identify Eq.˜53 as $\mathrm{p}({\bm{\mu}}^{\prime\prime}\mid{\bm{\mu}},{\bm{x}},\alpha+\alpha^{\prime})$. ∎
@@
on $\log\mathrm{p}({\bm{x}})$ (Kingma & Welling, 2013). In particular, we can choose our sequence of beliefs as the latent variable ${\bm{z}}=\{{\bm{\mu}}_{0},\ldots,{\bm{\mu}}_{k}\}$ and define the likelihood of $\bm{x}$} under this latent variable as
@@
If we now plug Eq.˜62 into the first term of Eq.˜60, we get
@@
which equals the definition of $\mathcal{L}_{\mathrm{R}}$ after plugging in Eq.˜61.
Next, we investigate the KL-divergence in Eq.˜60. We begin by defining the latent prior $\mathrm{p}({\bm{z}})$ autoregressively as
@@
Now, we plug Eqs.˜62 and 68 into the KL-divergence term from Eq.˜60.
@@
For $\mathcal{L}_{\mathrm{M}}^{k}$, we will consider its sum form from Eq.˜70.
@@
Note that $h(\lambda_{i-1})$ is a deterministic function of $\lambda_{i-1}$ and $\lambda_{0},\ldots,\lambda_{k}$ is a partition of the interval $\lambda_{0},\lambda_{0}+\alpha_{m}]=[\lambda_{0},\lambda_{\mathrm{M}}]$ with interval lengths of $\alpha_{i}$. It follows that Eq.˜71 is a Riemann sum. Since $f_{{\bm{\theta}}}$ is a neural network, we can assume that $h(\lambda_{i-1})$ is continuous almost everywhere. Combined with the fact that the interval lengths $\{\alpha_{i}\}$ converge uniformly to 0, it follows that $\mathcal{L}_{\mathrm{M}}^{k}$ converges to the Riemann integral
@@
We know from Eq.˜72 that $\mathcal{L}_{\mathrm{M}}^{\infty}$ is the following Riemann integral.

### v3

@@ 6.1 ImageNet
ImageNet64 ($1\text{\,}\mathrm{M}$ train steps)
BFN $3.222(6)$ $38.2(8)$
VDM $3.228(6)$ $35.2(7)$
BSI $3.218(4)$ $30.1(4)$
For this evaluation, we train a DiT (Peebles and Xie, 2023) in the BFN, VDM and BSI model, respectively, on the official 32×32 and 64×64 versions of ImageNet (Chrabaszcz et al., 2017). We train each model from three seeds and evaluate the log-likelihood of the test set in bits per dimension (BPD) and the sample quality in terms of Fréchet inception distance (FID) against the test set. For the log-likelihood, we evaluate each model’s ELBO with $5$ samples from the respective equivalent of $\mathcal{L}_{\mathrm{M}}^{\infty}$ and $2$ samples from the respective equivalent of $\mathcal{L}_{\mathrm{R}}$. For the sample quality, we draw 50 00050\,000 unconditional samples from each model with $1024$ steps and then compute the FID between the generated samples and the test set. On the 32×32 resolution images, we train the DiT-L-2 configuration for $2\text{\,}\mathrm{M}$ steps and on the 64×64 resolution data, we train the DiT-L-4 configuration for $1\text{\,}\mathrm{M}$ steps.
Table˜2 shows that BSI achieves equivalent likelihoods to VDM and BFN while generating higher-quality samples in terms of FID. This aligns with the result for consistency models by Song and Dhariwal (2024) that a larger variance of the initial state – initial belief ${\bm{\mu}}_{0}$ for BSI – improves the sample diversity. Ordering the models by improving FID, we have BFN first with an initial variance of 0 (${\bm{\mu}}_{0}={\bm{0}}$), then VDM with an initial variance of $1$ and finally BSI with an initial variance of $\lambda_{0}^{\shortminus{}\hskip-0.61455pt1}=100$. The magnitude of the FID on ImageNet64 aligns with the results reported by Peebles and Xie (2023) after $1\text{\,}\mathrm{M}$ training steps. Furthermore, Fig.˜3 shows that BSI achieves better sample quality than BFN and VDM for the same number of steps on ImageNet64.
@@
Figure 6: As $\lambda$ increases,
$\hat{{\bm{x}}}=f_{{\bm{\theta}}}({\bm{\mu}},\lambda)$ and the belief $\bm{\mu}$} converge.
Figure 7: $\lambda_{0}$ balances likelihood and sample quality for varying sample steps $k$.
Figure 8: FID and likelihood difference between models trained on $\mathcal{L}_{\mathrm{M}}^{k}$ and $\mathcal{L}_{\mathrm{M}}^{\infty}$ when evaluated for $k^{\prime}$ steps.
@@ 6.2 CIFAR10
We train the same U-Net architecture as VDM (Kingma et al., 2023) and BFN (Graves et al., 2023) on CIFAR10. Table˜4 shows that BSI achieves equivalent log-likelihoods in terms of BPD. Due to the significant number of training steps ($10\text{\,}\mathrm{M}$), we followed (Kingma et al., 2023; Graves et al., 2023) and trained only a single model on this dataset.
@@
6.3 Parameter Studies
In the following, we evaluate the impact of our modeling and parameter choices. Unless otherwise stated, we trained each model for $100\text{\,}\mathrm{k}$ steps on ImageNet32 with a DiT architecture, evaluated the likelihood of the test set in BPD with the infinite-step ELBO and used $1024$ sampling steps to compute the FID. We will verify the assumptions of the log-uniform proposal distribution $\mathrm{p}(\lambda)$, compare DiT and U-Net model architectures, and evaluate the prior precision $\lambda_{0}$ and training on the finite-step ELBO $\mathcal{L}_{\mathrm{M}}^{k}$.
Proposal Distribution
[…]
Table 3: Trained with U-Net architecture.
Model BPD FID
BFN $3.505(1)$ $14.2(4)$
VDM $3.527(9)$ $15.4(15)$
BSI $3.510(9)$ $12.8(6)$
Model Architecture
To ensure that the improvements in sample quality on ImageNet arise from BSI as a method and not from the architecture of the underlying model, we have also trained U-Nets on ImageNet32. Table˜3 shows that the U-Net exhibits the same characteristics as the DiT that we trained in Section˜6.1, i.e. equivalent likelihoods between BFN, VDM and BSI with a consistent improvement in FID. We chose the U-Net parameterization of (Kingma et al., 2023), which is also listed in Appendix˜E.
Initial Precision
In Fig.˜7, we evaluate the impact of the initial precision $\lambda_{0}$ on likelihood and sample quality. While the likelihood of test data improves with falling $\lambda_{0}$, i.e. increasing initial noise, the sample quality depends on the number of sampling steps. For a large number of steps, larger $\lambda_{0}$ perform slightly better, but with fewer steps an intermediate $\lambda_{0}$ is preferred. With fewer total sampling steps, decreasing $\lambda_{0}$ ensures that the sampling process still spends enough steps in the intermediate noise range, which is responsible for the generation of large-scale features in the images (Rissanen et al., 2022).
Training with $\mathcal{L}_{\mathrm{M}}^{\infty}$
By default, we train by optimizing the measurement loss $\mathcal{L}_{\mathrm{M}}^{\infty}$ of the infinite-step ELBO, but in practice the model will only use finitely many steps. Fig.˜8 shows that training on $\mathcal{L}_{\mathrm{M}}^{k}$ does not confer a consistent advantage in sample quality or likelihood. This justifies training by optimizing $\mathcal{L}_{\mathrm{M}}^{\infty}$ regardless of the number of steps used later and eliminates $k$ as a hyperparameter.
@@ 7 Conclusion
Software
For our results, we rely on excellent software packages, notably numpy (Harris et al., 2020), pytorch (Paszke et al., 2019), einops (Rogozhnikov, 2022), matplotlib (Hunter, 2007), h5py (Collette, 2013), hydra (Yadan, 2019) and jupyter (Granger and Pérez, 2021).
@@ Acknowledgments
This research was funded by the Bavarian State Ministry for Science and the Arts within the framework of the Geothermal Alliance Bavaria project. We want to thank the Munich Center for Machine Learning for providing compute resources and are thankful to David Lüdke and Jonas Dornbusch for their valuable feedback.
@@ A.2 Diffusion Models
While DMs initially achieved prominence in image generation (Dhariwal and Nichol, 2021), they have since been applied successfully across a variety of domains, such as text-to-image mapping (Saharia et al., 2022), fluid simulations (Lienen et al., 2024; Saydemir et al., 2024), adversarial attacks (Kollovieh et al., 2024a), temporal (Lüdke et al., 2023) and general point processes (Lüdke et al., 2024), molecular dynamics (Lewis et al., 2025), molecular structure generation (Ayadi et al., 2024), and time series forecasting (Kollovieh et al., 2023, 2024b).
@@
As we have shown in Section˜A.1, BFN are a special case of our generative framework in Section˜3 if we choose $\gamma_{0}=\infty$. Curiously, Eq.˜24 shows that this is the only prior on ${\bm{\mu}}_{0}$ for which the associated forward process is Markov as the coefficient of $\bm{x}$} becomes 0. This agrees with Xue et al. (2024), who have shown that BFN admit a formulation based on stochastic differential equations (SDEs), like score-based DMs.
@@
A.3 Stochastic Interpolants
Stochastic interpolants are a broad class of continuous-time stochastic processes that can interpolate between any two probability distributions $\rho_{0}$ and $\rho_{1}$ (Albergo et al., 2025). They also prescribe how to learn the interpolants’ dynamics to construct generative models and it is instructive to see how they relate to BSI. The subclass of spatially linear one-sided interpolants assumes that $\rho_{0}$ is a standard Normal distribution and defines the interpolant
[…]
where ${\bm{x}}_{1}\sim\rho_{1}$ and ${\bm{z}}_{t}\sim\mathcal{N}({\bm{0}},{\bm{I}})$. Given that $\alpha$ and $\beta$ are smooth, non-negative functions with $\alpha(0)=\beta(1)=1$ and $\alpha(1)=\beta(0)=0$, ${\bm{x}}_{t}$ smoothly interpolates between a standard Normal and the data distribution $\rho_{1}$.
For BSI, we can interpret the belief mean ${\bm{\mu}}_{\lambda}$ for a data sample $\bm{x}$}
${\bm{\mu}}_{\lambda}=\frac{1}{\sqrt{\lambda}}{\bm{z}}+\frac{\lambda-\lambda_{0}}{\lambda}{\bm{x}}$ (33)
as an interpolant that is normally distributed with mean $\bm{0}$} and precision $\lambda_{0}$ at $\lambda=\lambda_{0}$ and equals the data sample $\bm{x}$} at $\lambda=\infty$. We could rewrite this as an interpolant in the above sense on $[0,1]$ by parameterizing $\lambda$ as a strictly increasing function $\lambda(t):[0,1]\to\mathbb{R}_{+}$ with $\lambda(0)=\lambda_{0}$ and $\lim_{t\to 1}\lambda(t)=\infty$, e.g. $\lambda(t)=\lambda_{0}-\log(1-t)$, similar to the mapping between score-based diffusion and stochastic interpolants (Albergo et al., 2025, Section 5.1). But to avoid the scaling and correction factors, we will consider it an interpolant on $\lambda_{0},\infty]$ instead with $\alpha(\lambda)=1/\sqrt{\lambda}$ and $\beta(\lambda)=(\lambda-\lambda_{0})/\lambda$. We will furthermore write $\alpha(\lambda)$ and $\beta(\lambda)$ as $\alpha$ and $\beta$ to reduce visual clutter.
(Albergo et al., 2025, Section 4.4) shows that the probability path $\rho(\lambda,{\bm{\mu}}_{\lambda})$ of the interpolant solves the transport equation $\partial_{\lambda}\rho+\nabla\cdot(b\rho)=0$ with the velocity field
$b(\lambda,{\bm{\mu}})=\frac{\dot{\alpha}}{\alpha}{\bm{\mu}}+\bigg(\dot{\beta}-\frac{\dot{\alpha}}{\alpha}\beta\bigg)\eta(\lambda,{\bm{\mu}})$ (34)
and its score $\nabla\log\rho(\lambda,{\bm{\mu}})$ is given by
$s(\lambda,{\bm{\mu}})=-\frac{{\bm{\mu}}-\beta\eta(\lambda,{\bm{\mu}})}{\alpha^{2}}.$ (35)
$\eta(\lambda,{\bm{\mu}})=\operatorname*{\mathbb{E}}[{\bm{x}}\mid{\bm{\mu}}_{\lambda}={\bm{\mu}}]$ is the denoiser, i.e. the expected data sample that led to belief $\bm{\mu}$} at precision $\lambda$. Note that $\eta(\lambda,{\bm{\mu}})$ is learned with a model $\hat{\eta}$, which is equivalent to $f_{{\bm{\theta}}}({\bm{\mu}},\lambda)$ in BSI and fit by minimizing (Albergo et al., 2025, Eq. (4.21))
$\mathcal{L}(\hat{\eta})=\int_{\lambda_{0}}^{\lambda_{\mathrm{M}}}\operatorname*{\mathbb{E}}\bigg[\frac{1}{2}{\|\hat{\eta}(\lambda,{\bm{\mu}})\|}_{2}^{2}-{\bm{\mu}}\cdot\hat{\eta}(\lambda,{\bm{\mu}})\bigg]\,\mathrm{d}\lambda.$ (36)
[…]
With the velocity field and score, we can write down the forward SDE corresponding to the probability path $\rho$ as (Albergo et al., 2025, Corollary 18)
$\mathrm{d}{\bm{\mu}}_{\lambda}=b_{\mathrm{F}}(\lambda,{\bm{\mu}}_{\lambda})\,\mathrm{d}\lambda+\sqrt{2\varepsilon(\lambda)}\,\mathrm{d}W_{\lambda}$ (37)
where $W_{\lambda}$ is Brownian motion, $\varepsilon(\lambda):\mathbb{R}\to\mathbb{R}_{+}$ is any noise level specification and
$b_{\mathrm{F}}(\lambda,{\bm{\mu}})$ $=b(\lambda,{\bm{\mu}})+\varepsilon(\lambda)s(\lambda,{\bm{\mu}})$ (38)
$=\bigg(\frac{\dot{\alpha}}{\alpha}-\frac{\varepsilon}{\alpha^{2}}\bigg){\bm{\mu}}+\bigg(\frac{\varepsilon}{\alpha^{2}}-\frac{\dot{\alpha}}{\alpha}+\frac{\dot{\beta}}{\beta}\bigg)\beta\eta(\lambda,{\bm{\mu}})$
is the forward drift. If we plug in $\alpha$ and $\beta$, we get
$\mathrm{d}{\bm{\mu}}_{\lambda}=\bigg[-\bigg(\frac{1}{2\lambda}+\varepsilon\lambda\bigg){\bm{\mu}}_{\lambda}+\bigg(\varepsilon\lambda+\frac{1}{2\lambda}+\frac{\lambda_{0}}{\lambda(\lambda-\lambda_{0})}\bigg)\frac{\lambda-\lambda_{0}}{\lambda}\eta(\lambda,{\bm{\mu}}_{\lambda})\bigg]\,\mathrm{d}\lambda+\sqrt{2\varepsilon}\,\mathrm{d}W_{\lambda}.$ (39)
Since this holds for any non-negative $\varepsilon$, we can choose $\varepsilon=\frac{1}{2\lambda^{2}}$ to simplify the equation to
[…]
We can now sample from the learned stochastic interpolant by integrating Eq.˜40 from $\lambda_{0}$ to $\lambda_{M}$ (Albergo et al., 2025, Algorithm 5). Let’s say we are at precision $\lambda$ with state ${\bm{\mu}}_{\lambda}$ and want to move ahead by a step of length $\alpha$. With the Euler-Maruyama method suggested by (Albergo et al., 2025), the integration step becomes
${\bm{\mu}}_{\lambda+\alpha}$ $={\bm{\mu}}_{\lambda}+\alpha\cdot\frac{1}{\lambda}\big[\hat{\eta}(\lambda,{\bm{\mu}}_{\lambda})-{\bm{\mu}}_{\lambda}\big]+\frac{\sqrt{\alpha}}{\lambda}\,\bm{\varepsilon}$ (41)
$=\frac{\lambda-\alpha}{\lambda}{\bm{\mu}}_{\lambda}+\frac{\alpha}{\lambda}\hat{\eta}(\lambda,{\bm{\mu}}_{\lambda})+\frac{\sqrt{\alpha}}{\lambda}\,\bm{\varepsilon}$
where $\bm{\varepsilon}\sim\mathcal{N}({\bm{0}},{\bm{I}})$. This is almost the same as the BSI sampling step
${\bm{\mu}}_{i}$ $=\frac{\lambda_{i-1}{\bm{\mu}}_{i-1}+\alpha_{i}\big(\hat{{\bm{x}}}_{i-1}+\sqrt{\frac{1}{\alpha_{i}}}\bm{\varepsilon}_{i}\big)}{\lambda_{i-1}+\alpha_{i}}$ (42)
$=\frac{\lambda_{i}-\alpha_{i}}{\lambda_{i}}{\bm{\mu}}_{i-1}+\frac{\alpha_{i}}{\lambda_{i}}\hat{{\bm{x}}}_{i-1}+\frac{\sqrt{\alpha_{i}}}{\lambda_{i}}\bm{\varepsilon}_{i}$
in ˜8. The subtle difference is that Eq.˜41 uses the current $\lambda$ to compute ${\bm{\mu}}_{\lambda+\alpha}$ whereas Eq.˜42 uses the next $\lambda_{i}=\lambda_{i-1}+\alpha_{i}$ for ${\bm{\mu}}_{i}$. This difference reflects that BSI employs an exact Bayesian update rather than a first-order Euler-Maruyama approximation.
In summary, we can write BSI as a spatially linear one-sided stochastic interpolant, though two important differences remain. First, while the sampling steps Eqs.˜41 and 42 are equivalent in the continuous limit of $\alpha\to 0$, they differ in practice due to their derivation from an SDE discretization and posterior inference, respectively. Second, stochastic interpolants require the interpolation’s endpoints to equal the noise and data distribution exactly, which corresponds to $\lambda=\infty$ in the above formulation. In contrast, BSI only infers the sample $\bm{x}$} up to a maximum precision of $\lambda_{\mathrm{M}}$, e.g. precisely enough to identify the exact color in an image with 8-bit color channels.
@@ Appendix C Preconditioning Derivation
up to constant factors only depending on $\lambda$. If we plug in Eq.˜47 and isolate $f_{{\bm{\theta}}}^{\prime}$, this distance becomes
@@
Finally, we can plug $c_{\mathrm{skip}}$ into Eq.˜55 to get
@@ Appendix D Proofs
we can identify Eq.˜64 as $\mathrm{p}({\bm{\mu}}^{\prime\prime}\mid{\bm{\mu}},{\bm{x}},\alpha+\alpha^{\prime})$. ∎
@@
on $\log\mathrm{p}({\bm{x}})$ (Kingma and Welling, 2013). In particular, we can choose our sequence of beliefs as the latent variable ${\bm{z}}=\{{\bm{\mu}}_{0},\ldots,{\bm{\mu}}_{k}\}$ and define the likelihood of $\bm{x}$} under this latent variable as
@@
If we now plug Eq.˜73 into the first term of Eq.˜71, we get
@@
which equals the definition of $\mathcal{L}_{\mathrm{R}}$ after plugging in Eq.˜72.
Next, we investigate the KL-divergence in Eq.˜71. We begin by defining the latent prior $\mathrm{p}({\bm{z}})$ autoregressively as
@@
Now, we plug Eqs.˜73 and 79 into the KL-divergence term from Eq.˜71.
@@
For $\mathcal{L}_{\mathrm{M}}^{k}$, we will consider its sum form from Eq.˜81.
@@
Note that $h(\lambda_{i-1})$ is a deterministic function of $\lambda_{i-1}$ and $\lambda_{0},\ldots,\lambda_{k}$ is a partition of the interval $\lambda_{0},\lambda_{0}+\alpha_{m}]=[\lambda_{0},\lambda_{\mathrm{M}}]$ with interval lengths of $\alpha_{i}$. It follows that Eq.˜82 is a Riemann sum. Since $f_{{\bm{\theta}}}$ is a neural network, we can assume that $h(\lambda_{i-1})$ is continuous almost everywhere. Combined with the fact that the interval lengths $\{\alpha_{i}\}$ converge uniformly to 0, it follows that $\mathcal{L}_{\mathrm{M}}^{k}$ converges to the Riemann integral
@@
We know from Eq.˜83 that $\mathcal{L}_{\mathrm{M}}^{\infty}$ is the following Riemann integral.